OpenStudy (anonymous):

$180 is to be split between you and your friend, with your friend receiving $15 less. how much will each of you receive?

OpenStudy (anonymous):

I assume the $180 is being split 50/50. In this case, let x=180, the amount to be split. For your friend, it would be 0.5x-15. For you, it would be 0.5x+15.

OpenStudy (gw2011):

This is the way I look at the problem. Let x = your share Then x-15 = friend's share x + (x-15) = 180 x + x - 15 = 180 2x - 15 = 180 2x = 195 x = $97.50 This is your share x - 15 = $97.50 -15 = $82.50 This is friend's share Therefore, you receive $97.50 and your friend receives $82.50